A medication is given at a dosage of 3.000 mg of medication per kg of body weight. If 0.1500 g of medication is given, then what was the patient’s weight in pounds (lbs)? There are 453.59g in 1 lb. Do not enter “lbs” as part of your answer. Give the answer in scientific notation.
1
Expert's answer
2019-01-29T05:02:37-0500

As 1g = 1000 mg, then 0.1500 g *1000 mg/g = 150.0 mg of medication was given.

Solve the proportion:

3.000 mg - 1 kg

150.0 mg - x kg

3.000/150.0 = 1/x

x= 50.00 kg

As there are 1000g in 1 kg, then the weight (in g) of the patient is 50.00 kg *1000 g/kg = 50000g

As there are 453.59g in 1 lb, then the weigh (in lb) of the patient is 50000g / 453.59 g/lb = 110. 232 lb = 1.102*102 (lb)

Answer: 1.102*102